40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10037     Entry time: Fri Jun 13 18:16:00 2014
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: Changes to the PD frequency response measurement system 

 

As we had planned yesterday (ELOG 10034) I and Eric Gustafson wanted to manually measure the transimpedence for REFL33. But on closer inspection I found the RF signal cable coming from the Photodiode REF DET (mounted on the POY table), that we were supposed to plug into the network analyzer, did not have an SMA connector at the end. There was just the Teflon and metal part sticking out of the insulation. So we disconnected the cable labeled REF DET from the PD and pulled it out to fix it. (POY table and from near the 1Y1 rack)

 

 

Being unable to locate any SMA male connectors in the 40m lab [pasternack PE4025], we headed over to Downs where Rich Abbott did a quick and awesome job of soldering the SMA connector and also teaching me in the process. I will write an ELOG on how to do a clean solder of the SMA connectors to the RF cable shortly for future reference.

 

 

Coming back to the 40m we rerouted the REF DET cable from near the 1Y1 rack and into the POY table. This job was done mostly by Eric. We were also unable to locate a torque wrench to tighten the cable at the PD’s end and had to leave it finger tight. Eric is planning to buy a new torque wrench as we will need it often.

 

 

Also, I cross checked the SwithList located at /users/alex.cole/switchList with the RF switch connections at 1Y1 rack and turns out it is consistent, except that at CH2 of the first switch where MC REFL was to be connected, there is a unlabeled cable. It might belong to the correct PD, but must be made sure of. The rest of the channels that are not mentioned in the list were unconnected on the RF switch.

 

Now instead of disconnecting REFL 33 to make measurements with the NA, we had to take out AS55 from the RF switch, as the former was very hard to remove without the torque wrench. Then Eric removed the optical fiber which was illuminating the AS55 (AS table) from its mount to hook it up to the power meter. But then we were not sure of how to operate the Laser diode controller (LDC 3744C) and decided to leave stuff as it is and continue either tomorrow or on Monday. Right now we closed the optical fiber of AS55 with a cap and it remains unmounted. The RF cables of REF DET and AS55 were left hanging near the 1Y1 rack.

ELOG V3.1.3-